# Principal stratification with recurrent events truncated by a terminal event: A nested Bayesian nonparametric approach

### v2

@@ 3.1.3 Random partition properties
When applied to recurrent event data with a terminal event, the standard DP induces a random partition that is predominantly determined by the recurrent events rather than the terminal event. We provide detailed discussions on this unfavorable property of the DP-induced partition in Web Appendix B.2. This section discusses how the random partition induced by our proposed prior addresses this limitation. Define $\mathcal{P}_{n}=(\mathcal{P}_{n,u},\mathcal{P}_{n,y}),$ where $\mathcal{P}_{n,u}=(s_{u,1},\dots,s_{u,n})$ and $\mathcal{P}_{n,y}=(s_{y,1},\dots,s_{y,n})$ denote random partitions defined by cluster allocation labels. Specifically, $s_{u,i}=j$ if $\phi_{i}=\phi^{*}_{j}$ is the $j$-th distinct value of the atom $\phi$, and $s_{y,i}=l$ if $\theta_{i}=\theta^{*}_{l\mid j}$ is the $l$-th distinct value of the atom $\theta$ within the $j$-th $\phi$-cluster. For each $\phi$-cluster $j$, let $\mathcal{S}_{j}=\{i:s_{u,i}=j\},$ so $n_{j}=|\mathcal{S}_{j}|$ is the number of individuals in the $j$-th $\phi$-cluster. Denote the total number of $\phi$-clusters by $M_{n}$. Within each $\phi$-cluster $j$, define $\mathcal{S}_{l\mid j}=\{i:s_{u,i}=j,s_{y,i}=l\}$, so $n_{l\mid j}=|\mathcal{S}_{l\mid j}|$ is the size of the $l$-th $\theta$-subcluster, and let $M_{n,j}$ be the number of $\theta$-subclusters within cluster $j$. The unique cluster-specific parameters are then $\phi^{*}=(\phi^{*}_{j})_{j=1}^{M_{n}}$ and $\theta^{*}=\quantity(\theta^{*}_{1\mid j},\dots,\theta^{*}_{M_{n,j}\mid j})_{j=1}^{M_{n}}$. Finally, define $\mathcal{P}_{n_{j},y}=(s_{y,i}:i\in\mathcal{S}_{j})$, $U_{j}^{*}=\{U_{i}:i\in\mathcal{S}_{j}\}$, $\overline{Y}_{j}^{*}=\{\overline{Y}_{i}:i\in\mathcal{S}_{j}\}$, $\overline{Y}_{l\mid j}^{*}=\{\overline{Y}_{i}:i\in\mathcal{S}_{l\mid j}\}$.
@@
$p(\mathcal{P}_{n},\phi^{*},\theta^{*}\mid\overline{Y}_{1:n},U_{1:n})\propto p(\mathcal{P}_{n})\prod_{j=1}^{M_{n}}p_{0\phi}(\phi_{j}^{*})\prod_{l=1}^{M_{n,j}}p_{0\theta}(\theta_{l\mid j}^{*})\prod_{j=1}^{M_{n}}\prod_{i\in\mathcal{S}_{j}}K\left(U_{i}\mid\overline{Y}_{i},\phi_{j}^{*}\right)\prod_{l=1}^{M_{n,j}}\prod_{i\in\mathcal{S}_{l\mid j}}K\left(\overline{Y}_{i}\mid\theta_{l\mid j}^{*}\right).$
@@
The prior on the random partition induced by the EDP is
Furthermore, the posterior of the random partition of the EDP model is
$p(\mathcal{P}_{n}\mid\overline{Y}_{1:n},U_{1:n})\propto\alpha_{\phi}^{M_{n}}\prod_{j=1}^{M_{n}}\int_{\Phi}\alpha_{\theta}(\phi)^{M_{n,j}}\frac{\Gamma(\alpha_{\theta}(\phi))\Gamma(n_{j})}{\Gamma(\alpha_{\theta}(\phi)+n_{j})}dP_{0\phi}(\phi)h_{u}\left(U_{j}^{*}\mid\overline{Y}_{j}^{*}\right)\prod_{l=1}^{M_{n,j}}\Gamma(n_{l\mid j})h_{y}\left(\overline{Y}_{l\mid j}^{*}\right),$ (9)
where $h_{u}\left(U_{j}^{*}\mid\overline{Y}_{j}^{*}\right)=\int_{\Phi}\prod_{i\in\mathcal{S}_{j}}K\left(U_{i}\mid\overline{Y}_{i},\phi\right)dP_{0\phi}(\phi)$ and $h_{y}\left(\overline{Y}_{l\mid j}^{*}\right)=\int_{\Theta}\prod_{i\in\mathcal{S}_{l\mid j}}K\left(\overline{Y}_{i}\mid\theta\right)dP_{0\theta}(\theta)$.
The proof is provided in Web Appendix C. Moreover, by marginalizing over all possible $\theta$-subpartitions for each $\phi$-cluster, one obtains the posterior for $\mathcal{P}_{n,u}$:
@@ 3.1.4 Posterior inference
$\begin{split}\mathcal{L}_{\mathrm{obs}}=&\prod_{i=1}^{n}f^{1-\delta^{C}_{i}}\left(U_{i}\mid Z_{i},\mathbf{X}_{i},\boldsymbol{\beta}_{u,i},\gamma_{i}^{Z_{i}},\tau_{i}\right)S_{U}^{\delta^{C}_{i}}\left(\log(\mathcal{T}_{i})\mid Z_{i},\mathbf{X}_{i},\boldsymbol{\beta}_{u,i},\gamma_{i}^{Z_{i}},\tau_{i}\right)\\ &\times\prod_{j=1}^{N_{i}}f\left(Y_{ij}\mid Z_{i},\mathbf{X}_{i},\mathbf{V}_{i(N_{i}+1)},\boldsymbol{\beta}_{y,ij},\psi_{ij},\gamma_{i}^{Z_{i}},\sigma_{ij}\right)\\ &\times\penalty 10000\ S_{Y}\left(\log(\mathcal{T}_{i}-T_{iN_{i}})\mid Z_{i},\mathbf{X}_{i},\mathbf{V}_{i(N_{i}+1)},\boldsymbol{\beta}_{y,i(N_{i}+1)},\psi_{i(N_{i}+1)},\gamma_{i}^{Z_{i}},\sigma_{i(N_{i}+1)}\right),\end{split}$
@@
where $Y^{*}$ and $U^{*}$ represent the imputed values of the gap time and survival time, respectively. This complete likelihood admits the standard posterior sampling technique based on the densities. Specifically, we employ an approximated blocked Gibbs sampler based on a two-level truncation of the square-breaking representation of the EDP proposed by Burns2023. In this algorithm, we first select conservative upper bounds on the number of latent classes of the square-breaking representation of (6) as follows.
@@
Let $G_{i}\in\{1,...,K\}$ and $H_{ij}\in\{1,...,L\}$ denote the latent cluster indicators for individual $i=1,\ldots,n$ and time $j=1,\ldots,N_{i}$. Our data augmentation algorithm further imputes these latent indicators in each iteration to facilitate the posterior updates of model parameters. Specifically, we specify Multinomial distributions $G_{i}\sim\mathrm{MN}(\mathbf{w}^{\phi})$ on $G_{i}$ and $H_{ij}\sim\mathrm{MN}(\mathbf{w}^{\theta}_{k})$ on $H_{ij}$, where $\mathbf{w}^{\phi}=(w^{\phi}_{1},\ldots,w^{\phi}_{K})^{\top}$ and $\mathbf{w}^{\theta}_{k}=(w^{\theta}_{1\mid k},\ldots,w^{\theta}_{L\mid k})^{\top}$ contains the weights from the EDDP. Burns2023 demonstrated that an accurate approximation to the exact EDP is obtained as long as the truncation bound is sufficiently large. To ensure this, we ran several MCMC iterations with different values of $K$ and $L$ and increased them if all clusters were occupied. Overall, our algorithm iterates between drawing from the conditional distributions of censored outcomes, latent cluster indicators, and model parameters.
@@
The essential Gibbs sampler is outlined as follows: (i) given all model parameters, $G_{i}$ and $H_{i}$, sample $Y^{*}_{i(N_{i}+1)}$ and $U^{*}_{i}$, (ii) given all model parameters, $Y^{*}_{i(N_{i}+1)}$ and $U^{*}_{i}$, sample $G_{i}$ and $H_{i}$, (iii) given $Y^{*}_{i(N_{i}+1)}$, $U^{*}_{i}$, $G_{i}$ and $H_{i}$, sample all model parameters, and (iv) compute the estimands based on Theorem 1. Specifically, when imputing $(N_{i}+1)$-th gap time $Y^{*}_{i(N_{i}+1)}$ in the first step, we draw from its conditional predictive distribution $p(Y\mid Y>\log(\mathcal{T}_{i}-T_{iN_{i}}),-)$, where “-” denotes the rest of relevant variables, as the final gap time is always censored at $\mathcal{T}_{i}-T_{iN_{i}}$, the actual gap time should be greater than the censored gap time. Similarily, for the unit with $\delta_{i}^{C}=1$, its survival time $U_{i}$ from the posterior predictive distribution, enabling us to determine if the unit belongs to $\mathcal{AS}(r)$. For the imputation of $U_{i}$ with $\delta_{i}^{C}=1$, we also draw from its conditional predictive distribution $p(U\mid U>\log(\mathcal{T}_{i}-D_{i}),-)$ as the survival time is censored and its actual survival time should be greater than the observed one. When computing esimands in the last step, e.g., $\mu^{z}\quantity(t;z)$, we approximate it with its empirical distribution. For example, the numerator of $\mu^{z}\quantity(t;z)$ in Theorem 1 is approximated by $\frac{1}{n}\sum_{i=1}^{n}\kappa_{t,r}(Z_{i},X_{i},\boldsymbol{\gamma}_{i})\eta_{t,r}(1,X_{i},\boldsymbol{\gamma}_{i})\eta_{t,r}(0,X_{i},\boldsymbol{\gamma}_{i})$. This is a conventional approximation in the Bayesian paradigm (Li_Ding2023). See Web Appendix D for full details of the algorithm. The initial parameter values were randomly drawn from the prior distributions, and the posterior samples were obtained by running a chain for $5000$ MCMC iterations after an initial $20000$ burn-in iteration. Convergence was monitored by the trace plots, confirming that the chains had reached stationarity with good mixing.
@@ 4 Simulation Studies
In this section we examine the performance of the proposed methods through simulation studies. Specifically, we evaluate the frequentist properties of the proposed EDDP prior for estimating the key estimands $\mu^{z}(t;r)$ and $\tau^{z}(t;r)$ for $z\in\{0,1\}$, with comparisons to methods commonly used in the recurrent event analysis: linear model (LM) (e.g., Comment2019; Lyu2023), Dirichlet process mixture (DPM) (e.g., Paulon2020; Tian2024), and dependent Dirichlet process mixture (DDPM) (e.g., Xu2022).
@@
We simulate $100$ datasets with $n=2000$ individuals and evaluate the bias and root mean square error (RMSE) of a point estimator (posterior mean), as well as the frequentist coverage probability (CP) and average length (AL) of the $95\%$ central credible interval. The dataset is generated from the following mixture models with random seeds. $U_{i}^{z}\sim\sum_{k=1}^{3}w_{k}\mathrm{N}(\alpha^{u}+\mathbf{X}_{i}^{\top}\phi_{k}+z\beta^{u}_{k}+\gamma_{i}^{z},0.2)$, $Y_{ij}^{z}\sim\sum_{k=1}^{3}\pi_{k}\mathrm{N}(\alpha^{y}+\mathbf{X}_{i}^{\top}\theta_{k}+z\beta^{y}_{k}+\psi\gamma_{i}^{z},0.2),$ where $\mathbf{X}_{i}\sim\mathrm{MVN}((0,0,0)^{\top},I_{3})$ with $I_{d}$ is the identity matrix of dimension $d$, $Z_{i}\sim\mathrm{Bern}(0.5)$, $(w_{1},w_{2},w_{3})=(0.3,0.4,0.3)$, $(\pi_{1},\pi_{2},\pi_{3})=(0.3,0.4,0.3)$, $\alpha^{u}=6.5$, $\alpha^{y}=5.0$, $\phi_{1}=(0.2,0.15,-0.1)^{\top}$, $\phi_{2}=0.5\phi_{1}$, $\phi_{3}=0.3\phi_{1}$, $\theta_{1}=(0.15,0.10,-0.05)^{\top}$, $\theta_{2}=0.5\theta_{1}$, $\theta_{3}=0.3\theta_{1}$, $(\beta^{u}_{1},\beta^{u}_{2},\beta^{u}_{3})=(-1.0,-0.5,-1.3)$, $(\beta^{y}_{1},\beta^{y}_{2},\beta^{y}_{3})=(-0.4,-0.2,-0.52)$, and $\psi=0.1$. The frailty follows the multivariate log-normal distribution such that $\boldsymbol{\gamma}_{i}=\exp(\boldsymbol{\gamma}_{i}^{\prime})$, where $\boldsymbol{\gamma}_{i}^{\prime}\sim\mathrm{MVN}((0,0)^{\top},\left(\begin{smallmatrix}0.2&0.2\rho\\ 0.2\rho&0.2\end{smallmatrix}\right))$ with a correlation parameter $\rho=0.5$. The mixture components $w_{k}$ and $\pi_{k}$ take the same weight value for $k=1,2,3$, but the assignment to each distribution component is independent. The observed survival time and recurrent gap times are obtained by transforming the log-scale variables $U_{i}^{z}$ and $Y_{ij}^{z}$ into $D_{i}=\exp(U_{i}^{Z_{i}})$ and $T_{ij}=\exp(Y_{ij}^{Z_{i}})$, and $T_{ij}$ is generated recurrently until the sum exceeds the censoring time $\mathcal{T}_{i}=\min(D_{i},C_{i})$ where $C_{i}\sim\mathrm{Unif}(300,1000)$, where $C_{i}$ represents the independent censoring time for each individual. The estimands’ true values are approximated by Monte Carlo simulation using large samples.
For inference, we use model (7) for the EDDPM. In the DDPM, the nested prior structure is removed from the EDDPM. In the DPM, covariate dependence is further removed from the DDPM, and the frailty and error terms are modeled using the DPM, resulting in a model equivalent to that of Paulon2020. For the LM, we fit a linear mixed model to the log‑transformed survival and gap times. The specification includes a subject‑level frailty term and assumes normal residual errors, which yields the familiar accelerated failure time (AFT) representation. This setup is close to the models explored by Comment2019; Lyu2023. The only difference is that their proportional hazard model assumes an extreme‑value distribution for the error term. We use proper, weakly informative conjugate priors for all parameters. Specifically, we choose the hyperparameters $a_{\alpha_{\phi}}=\alpha_{\theta\mid\phi}=a_{\tau}=a_{\sigma}=2.0$, $b_{\alpha_{\phi}}=b_{\alpha_{\theta\mid\phi}}=b_{\tau}=b_{\sigma}=1.0$, $\mu_{\gamma}=\mu_{\psi}=0.0$, $\sigma^{2}_{\gamma}=\sigma^{2}_{\psi}=3.0^{2}$, $\boldsymbol{\mu}_{\boldsymbol{\beta}_{u}}$ and $\boldsymbol{\mu}_{\boldsymbol{\beta}_{y}}$ are zero-vectors of appropriate sizes, and $\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{u}}$ and $\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{y}}$ are diagonal matrices with $3.0^{2}$ on the diagonal elements.
Table 2 presents the simulation results. Overall, the results consistently show that our methods have the smallest bias and RMSE across different scenarios, indicating EDDPM’s superior accuracy in point estimation. Turning to interval estimation, which is assessed by the coverage probability (CP) of $95\%$ credible intervals, EDDPM stands out for its well-calibrated CP, closer to the target $95\%$. The linear model is sensitive to the latent treatment heterogeneity of the data-generating processes, yielding the worst performance across most metrics. Although the DPM and DDPM models improve upon the linear model, their performance falls short of the exemplary standards set by EDDPM. In practice, we recommend evaluating the performance of these models, as illustrated in Section 5.

### v3

@@ 3.3.3 Random partition properties
When applied to recurrent event data with a terminal event, the standard DP induces a random partition that is predominantly determined by the recurrent events rather than the terminal event. We provide detailed discussions on this unfavorable property of the DP-induced partition in Supplementary Material Section A.2. This section discusses how the random partition induced by our proposed prior addresses this limitation. Define $\mathcal{P}_{n}=(\mathcal{P}_{n,u},\mathcal{P}_{n,y}),$ where $\mathcal{P}_{n,u}=(s_{u,1},\dots,s_{u,n})$ and $\mathcal{P}_{n,y}=(s_{y,1},\dots,s_{y,n})$ denote random partitions defined by cluster allocation labels. Specifically, $s_{u,i}=j$ if $\phi_{i}=\phi^{*}_{j}$ is the $j$-th distinct value of the atom $\phi$, and $s_{y,i}=l$ if $\theta_{i}=\theta^{*}_{l\mid j}$ is the $l$-th distinct value of the atom $\theta$ within the $j$-th $\phi$-cluster. For each $\phi$-cluster $j$, let $\mathcal{S}_{j}=\{i:s_{u,i}=j\},$ so $n_{j}=|\mathcal{S}_{j}|$ is the number of individuals in the $j$-th $\phi$-cluster. Denote the total number of $\phi$-clusters by $M_{n}$. Within each $\phi$-cluster $j$, define $\mathcal{S}_{l\mid j}=\{i:s_{u,i}=j,s_{y,i}=l\}$, so $n_{l\mid j}=|\mathcal{S}_{l\mid j}|$ is the size of the $l$-th $\theta$-subcluster, and let $M_{n,j}$ be the number of $\theta$-subclusters within cluster $j$. The unique cluster-specific parameters are then $\phi^{*}=(\phi^{*}_{j})_{j=1}^{M_{n}}$ and $\theta^{*}=\quantity(\theta^{*}_{1\mid j},\dots,\theta^{*}_{M_{n,j}\mid j})_{j=1}^{M_{n}}$. Finally, define $\mathcal{P}_{n_{j},y}=(s_{y,i}:i\in\mathcal{S}_{j})$, $U_{j}^{*}=\{U_{i}:i\in\mathcal{S}_{j}\}$, $\overline{Y}_{j}^{*}=\{\overline{Y}_{i}:i\in\mathcal{S}_{j}\}$, $\overline{Y}_{l\mid j}^{*}=\{\overline{Y}_{i}:i\in\mathcal{S}_{l\mid j}\}$.
@@
$\begin{split}p(\mathcal{P}_{n},\phi^{*},\theta^{*}&\mid\overline{Y}_{1:n},U_{1:n})\\ \propto&p(\mathcal{P}_{n})\prod_{j=1}^{M_{n}}p_{0\phi}(\phi_{j}^{*})\prod_{l=1}^{M_{n,j}}p_{0\theta}(\theta_{l\mid j}^{*})\prod_{j=1}^{M_{n}}\prod_{i\in\mathcal{S}_{j}}K\left(U_{i}\mid\overline{Y}_{i},\phi_{j}^{*}\right)\prod_{l=1}^{M_{n,j}}\prod_{i\in\mathcal{S}_{l\mid j}}K\left(\overline{Y}_{i}\mid\theta_{l\mid j}^{*}\right).\end{split}$
@@
Let $h_{u}\left(U_{j}^{*}\mid\overline{Y}_{j}^{*}\right)=\int_{\Phi}\prod_{i\in\mathcal{S}_{j}}K\left(U_{i}\mid\overline{Y}_{i},\phi\right)dP_{0\phi}(\phi)$ and $h_{y}\left(\overline{Y}_{l\mid j}^{*}\right)=\int_{\Theta}\prod_{i\in\mathcal{S}_{l\mid j}}K\left(\overline{Y}_{i}\mid\theta\right)dP_{0\theta}(\theta)$. The posterior of the random partition of the EDP model is
$\begin{split}&p(\mathcal{P}_{n}\mid\overline{Y}_{1:n},U_{1:n})\\ &\propto\alpha_{\phi}^{M_{n}}\prod_{j=1}^{M_{n}}\int_{\Phi}\alpha_{\theta}(\phi)^{M_{n,j}}\frac{\Gamma(\alpha_{\theta}(\phi))\Gamma(n_{j})}{\Gamma(\alpha_{\theta}(\phi)+n_{j})}dP_{0\phi}(\phi)h_{u}\left(U_{j}^{*}\mid\overline{Y}_{j}^{*}\right)\prod_{l=1}^{M_{n,j}}\Gamma(n_{l\mid j})h_{y}\left(\overline{Y}_{l\mid j}^{*}\right).\end{split}$ (7)
The proof is provided in Supplementary Material Section B. Moreover, by marginalizing over all possible $\theta$-subpartitions for each $\phi$-cluster, one obtains the posterior for $\mathcal{P}_{n,u}$:
@@ 3.3.4 Posterior inference
$\begin{split}\mathcal{L}_{\mathrm{obs}}=&\prod_{i=1}^{n}f^{1-\delta^{C}_{i}}\left(U_{i}\mid Z_{i},\mathbf{X}_{i},\boldsymbol{\beta}_{u,i},\gamma_{i}^{Z_{i}},\tau_{i}\right)S_{U}^{\delta^{C}_{i}}\left(\log(\mathcal{T}_{i})\mid Z_{i},\mathbf{X}_{i},\boldsymbol{\beta}_{u,i},\gamma_{i}^{Z_{i}},\tau_{i}\right)\\ &\times\prod_{j=1}^{N_{i}}f\left(Y_{ij}\mid Z_{i},\mathbf{X}_{i},\mathbf{V}_{ij},\boldsymbol{\beta}_{y,ij},\psi_{ij},\gamma_{i}^{Z_{i}},\sigma_{ij}\right)\\ &\times\penalty 10000\ S_{Y}\left(\log(\mathcal{T}_{i}-T_{iN_{i}})\mid Z_{i},\mathbf{X}_{i},\mathbf{V}_{i(N_{i}+1)},\boldsymbol{\beta}_{y,i(N_{i}+1)},\psi_{i(N_{i}+1)},\gamma_{i}^{Z_{i}},\sigma_{i(N_{i}+1)}\right),\end{split}$
@@
where $Y^{*}$ and $U^{*}$ represent the imputed values of the gap time and survival time, respectively. This complete likelihood admits the standard posterior sampling technique based on the densities. Specifically, we employ an approximate blocked Gibbs sampler based on a two-level truncation of the square-breaking representation of the EDP proposed by Burns and Daniels (2023). In this algorithm, we first select conservative upper bounds on the number of latent classes of the square-breaking representation of (4) as follows.
@@
Let $G_{i}\in\{1,...,K\}$ and $H_{ij}\in\{1,...,L\}$ denote the latent cluster indicators for individual $i=1,\ldots,n$ and time $j=1,\ldots,N_{i}$. Our data augmentation algorithm further imputes these latent indicators in each iteration to facilitate the posterior updates of model parameters. Specifically, we specify Multinomial distributions $G_{i}\sim\mathrm{MN}(\mathbf{w}^{\phi})$ on $G_{i}$ and $H_{ij}\sim\mathrm{MN}(\mathbf{w}^{\theta}_{k})$ on $H_{ij}$, where $\mathbf{w}^{\phi}=(w^{\phi}_{1},\ldots,w^{\phi}_{K})^{\top}$ and $\mathbf{w}^{\theta}_{k}=(w^{\theta}_{1\mid k},\ldots,w^{\theta}_{L\mid k})^{\top}$ contains the weights from the EDDP. Burns and Daniels (2023) demonstrated that an accurate approximation to the exact EDP is obtained as long as the truncation bound is sufficiently large. To ensure this, we ran several MCMC iterations with different values of $K$ and $L$ and increased them if all clusters were occupied. Overall, our algorithm iterates between drawing from the conditional distributions of censored outcomes, latent cluster indicators, and model parameters.
@@
The essential Gibbs sampler is outlined as follows: (i) given all model parameters, $G_{i}$ and $H_{i}$, sample $Y^{*}_{i(N_{i}+1)}$ and $U^{*}_{i}$, (ii) given all model parameters, $Y^{*}_{i(N_{i}+1)}$ and $U^{*}_{i}$, sample $G_{i}$ and $H_{i}$, (iii) given $Y^{*}_{i(N_{i}+1)}$, $U^{*}_{i}$, $G_{i}$ and $H_{i}$, sample all model parameters, and (iv) compute the estimands based on Theorem 1. Specifically, when imputing $(N_{i}+1)$-th gap time $Y^{*}_{i(N_{i}+1)}$ in the first step, we draw from its conditional predictive distribution $p(Y\mid Y>\log(\mathcal{T}_{i}-T_{iN_{i}}),\bullet)$, where $\bullet$” denotes the rest of relevant variables, as the final gap time is always censored at $\mathcal{T}_{i}-T_{iN_{i}}$, the actual gap time should be greater than the censored gap time. Similarly, for the unit with $\delta_{i}^{C}=1$, its survival time $U_{i}$ from the posterior predictive distribution enables us to determine if the unit belongs to $\mathcal{AS}(r)$. For the imputation of $U_{i}$ with $\delta_{i}^{C}=1$, we also draw from its conditional predictive distribution $p(U\mid U>\log(\mathcal{T}_{i}),\bullet)$ because the observed survival time is right-censored and the realized $U_{i}$ must exceed the censoring bound. When computing the estimands in step (iv), we approximate the integrals in (3) by Monte Carlo simulation. For example, for a fixed arm $z\in\{0,1\}$, the numerator of (3) for $\mu^{z}(t;r)$ is approximated by $\frac{1}{n}\sum_{i=1}^{n}\kappa_{t,r}\!\left(z,X_{i},\boldsymbol{\gamma}_{i}\right)\,\eta_{r}\!\left(1,X_{i},\boldsymbol{\gamma}_{i}\right)\,\eta_{r}\!\left(0,X_{i},\boldsymbol{\gamma}_{i}\right),$ with an analogous sample-average approximation for the denominator without the need to additionally model the distribution of baseline covariates. This is a conventional approximation in the Bayesian paradigm (Li et al., 2023). See Supplementary Material Section C for full details of the algorithm. The initial parameter values were randomly drawn from the prior distributions, and the posterior samples were obtained by running a chain for $10000$ MCMC iterations after an initial $40000$ burn-in iterations. Convergence was monitored by the trace plots, confirming that the chains had reached stationarity with good mixing.
@@ 4 Simulation Studies
We examine the performance of the proposed methods through simulation studies. Specifically, here we evaluate the frequentist properties of the proposed EDDP prior for estimating the key estimands $\mu^{z}(t;r)$ for $z\in\{0,1\}$. We simulate $300$ datasets with $n=1000$ individuals and report the bias and root mean square error (RMSE) of a point estimator (posterior mean), as well as the frequentist coverage probability (CP) and average length (AL) of the $95\%$ central credible interval. The dataset is generated from the following mixture models with random seeds:
$U_{i}^{z}$ $\sim\sum_{k=1}^{3}w_{k}\mathrm{N}(\alpha^{u}+\mathbf{X}_{i}^{\top}\phi_{k}+z\beta^{u}_{k}+\gamma_{i}^{z},0.2),$
$Y_{ij}^{z}$ $\sim\sum_{k=1}^{3}\pi_{k}\mathrm{N}(\alpha^{y}+\mathbf{X}_{i}^{\top}\theta_{k}+z\beta^{y}_{k}+\psi\gamma_{i}^{z},0.2),$
@@
where $\mathbf{X}_{i}\sim\mathrm{MVN}((0,0,0)^{\top},I_{3})$ with $I_{d}$ is the identity matrix of dimension $d$, $Z_{i}\sim\mathrm{Bern}(0.5)$, $(w_{1},w_{2},w_{3})=(0.3,0.4,0.3)$, $(\pi_{1},\pi_{2},\pi_{3})=(0.3,0.4,0.3)$, $\alpha^{u}=6.5$, $\alpha^{y}=5.0$, $\phi_{1}=(0.2,0.15,-0.1)^{\top}$, $\phi_{2}=-0.5\phi_{1}$, $\phi_{3}=0.3\phi_{1}$, $\theta_{1}=(0.25,-0.10,-0.15)^{\top}$, $\theta_{2}=-0.5\theta_{1}$, $\theta_{3}=0.3\theta_{1}$, $(\beta^{u}_{1},\beta^{u}_{2},\beta^{u}_{3})=(1.0,0.5,1.3)$, $(\beta^{y}_{1},\beta^{y}_{2},\beta^{y}_{3})=(0.5,0.25,0.65)$, and $\psi=0.1$. The frailty follows the multivariate log-normal distribution such that $\boldsymbol{\gamma}_{i}=\exp(\boldsymbol{\gamma}_{i}^{\prime})$, where $\boldsymbol{\gamma}_{i}^{\prime}\sim\mathrm{MVN}((0,0)^{\top},\left(\begin{smallmatrix}0.2&0.2\rho\\ 0.2\rho&0.2\end{smallmatrix}\right))$ with a correlation parameter $\rho=0.5$. The mixture components $w_{k}$ and $\pi_{k}$ take the same weight value for $k=1,2,3$, but the assignment to each distribution component is independent. The observed survival time and recurrent gap times are obtained by transforming the log-scale variables $U_{i}^{z}$ and $Y_{ij}^{z}$ into $D_{i}=\exp(U_{i}^{Z_{i}})$ and $W_{ij}=\exp(Y_{ij}^{Z_{i}})$, and $W_{ij}$ is generated recurrently until the sum exceeds the censoring time $\mathcal{T}_{i}=\min(D_{i},C_{i})$ where $C_{i}\sim\mathrm{Unif}(300,1000)$, where $C_{i}$ represents the independent censoring time for each individual. The estimands’ true values are approximated by Monte Carlo simulation using large samples.